# CHESS : Optimizing LLM Inference via Channel-Wise Thresholding and Selective Sparsification

### v1

@@ Abstract
Deploying large language models (LLMs) on edge devices presents significant challenges due to the substantial computational overhead and memory requirements. Activation sparsification can mitigate these challenges by reducing the number of activated neurons during inference. Existing methods typically employ thresholding-based sparsification based on the statistics of activation tensors. However, these methods do not explicitly model the impact of activation sparsification on performance, leading to suboptimal performance degradation. To address this issue, this paper reformulates the activation sparsification problem by introducing a new objective that optimizes the sparsification decisions. Building on this reformulation, we propose CHESS, a general activation sparsification approach via CHannel-wise thrEsholding and Selective Sparsification. First, channel-wise thresholding assigns a unique threshold to each activation channel in the feed-forward network (FFN) layers. Then, selective sparsification involves applying thresholding-based activation sparsification to specific layers within the attention modules. Finally, we detail the implementation of sparse kernels to accelerate LLM inference. Experimental results demonstrate that the proposed CHESS achieves lower performance degradation over 8 downstream tasks while activating fewer parameters compared to existing methods, thus speeding up the LLM inference by up to 1.27x.
@@
Junhui He1  Shangyu Wu2,3  Weidong Wen1  Chun Jason Xue2  Qingan Li1 1 Wuhan University  2 MBZUAI  3 City University of Hong Kong
@@ 1 Introduction
Large Language Models (LLMs) have prevailed in a wide range of applications across various fields, such as code generation tools, office assistants, input method editors, voice assistants, and assistive applications designed for individuals with disabilities. However, due to the substantial computation and memory requirements of LLM inferences, deploying LLMs on edge devices is still challenging. To mitigate these overheads, utilizing the inherent activation sparsity of LLM has emerged as a promising strategy (Liu et al., 2023; Song et al., 2023; Alizadeh et al., 2023). This approach has proven effective for models with the ReLU activation function (Li et al., 2023; Liu et al., 2023).
Contemporary LLMs demonstrate that SwiGLU or GeGLU activation functions can further boost the model performance, but they induce less activation sparsity. Consequently, several methods (Mirzadeh et al., 2024; Song et al., 2024) are proposed to explore more sparsity by regularizing the SwiGLU or GeGLU activation. However, those works require fine-tuning the LLMs, which entails significant training overhead. To avoid training overheads and improve activation sparsification in modern LLMs, Lee et al. (2024) propose a thresholding-based pruning method to actively sparsify the activation tensors during the inference stage. However, this thresholding technique focuses solely on the statistics of the activation tensors themselves, failing to model the impact of sparsification on overall model performance. This lack of modeling results in suboptimal performance degradation.
To address the above limitations, this paper proposes CHESS , a new activation sparsification optimization via CHannel-wise thrEsholding and Selective Sparsification. First, this paper introduces a new objective that models the impact of sparsification on model performance and optimizes the sparsification decisions. Then, this paper proposes channel-wise thresholding for FFN layers in LLMs, which determines the unique threshold for each activation channel. Furthermore, this paper proposes selective sparsification, which applies thresholding-based activation sparsification to the target submodules in the attention module. Finally, this paper presents the implementations of sparse kernels to accelerate the inference based on the sparse activations.
To validate the effectiveness of the proposed CHESS , this paper conducts comprehensive experiments on various downstream tasks and state-of-the-art LLMs. Experimental results demonstrate that the proposed CHESS can achieve a lower performance degradation while a better end-to-end inference speedup. Codes are available in  [^1] https://anonymous.4open.science/r/CHESS-BA40.
@@
  • This paper systematically formulates the activation sparsification problem and establishes a connection between sparsity and model performance.
  • This paper proposes two activation sparsifications, the channel-wise thresholding for FFN modules and the selective sparsification for Attention modules, which can be widely applied in existing LLMs.
  • To make full use of the activation sparsity, this paper presents the detailed algorithms for implementing the sparse kernels.
  • Experimental results demonstrate the efficacy and scalability of the proposed CHESS .
@@ 2.1 Activation Sparsification
Activation functions introduce non-linearity into neural networks, allowing networks to capture complex patterns in the data. ReLU (Glorot et al., 2011), as a popular activation function, has been widely applied in most neural networks for addressing gradient vanish issues (Zhang et al., 2022). Another benefit of ReLU is introducing the sparsity into the activation tensors. Recent studies (Li et al., 2023; Liu et al., 2023) have demonstrated this effect, showing that up to 95% of the intermediate FFN activations in OPT models are zero. Such sparsity can be used to accelerate the model inference while maintaining comparable model performance (Liu et al., 2023; Alizadeh et al., 2023; Song et al., 2023).
@@
To improve the activation sparsification in modern LLMs, existing work (Lee et al., 2024) proposes a thresholding-based pruning method called CATS on some activation tensors in FFN layers. CATS first computes the cutoff threshold over a subset of training data according to the given sparsity level, then sparsifies the activations during inference and achieves end-to-end speedup via efficient sparse kernel design. Although CATS can improve activation sparsification, it only focuses on the statistics of the activation tensors themselves without modeling the impact of activation sparsification on the model performance, leading to suboptimal performance drop.
@@ 2.2 Motivation
Following the observations in CATS (Lee et al., 2024), this paper also aims to apply activation sparsification in the Gated-MLP blocks of FFN modules. The formal expression of the gated-MLP block is defined as,
@@
Since the activation function introduces sparsity where the values of many elements in the output tensor are close to zero, we focus on pruning the output of the gate projection layer, i.e., $a^{\text{gate}}$. Then, the following computations, such as the matrix multiplication for $a^{\text{up}}$, the element-wise multiplication between $a^{\text{up}}$ and $a^{\text{gate}}$, or the matrix multiplication with $W^{\text{down}}$, can further be reduced due to the zero elements in the pruned $a^{\text{gate}}$.
@@
We decompose all activations in the pruned tensor into two subsets, i.e., the pruned $\hat{a}^{\text{gate}}_{\mathcal{P}}$ which are all zeros and the non-pruned $\hat{a}^{\text{gate}}_{\mathbb{U}-\mathcal{P}}$ which are the same as the corresponding values in $a^{\text{gate}}$. Thus, we can simplify the objective defined in Equation 3 as: finding a subset of indices $\mathcal{P}$ that indicates the index of the pruned elements, and satisfies sparsity level $|\mathcal{P}|\geq k\cdot|\mathbb{U}|$, while minimizing the sparsification error illustrated in Equation 4, where $\mathbb{U}=\{1,\ldots,d\}$, $d$ is the feature dimension of $a^{\text{gate}}$.
@@ 3 CHESS : Activation Sparsification via Channel-Wise Thresholding and Selective Sparsification
Figure 1: Distribution of absolute activation values $|a^{\text{up}}_{i}|$ across different inputs for various channels in the FFN of layer 16 in the Llama-3-8B model. Subplots (a), (b), and (c) show the distributions for channels 1241, 8718, and 12005 respectively. These plots illustrate that while the absolute activation values within a single channel remain relatively consistent across different inputs, there is significant variation between different channels.
@@
Figure 2: Distribution of $\|W_{i,:}\|_{2}^{2}$ of different rows $i$ in attention projections of layer 16 of Llama-3-8B. Subplots (a), (b) and (c) show the distributions for weights of query, key, value and output projection, respectively. These plots illustrate that all rows from the same weight tensor exhibit similar $\|W_{i,:}\|_{2}^{2}$.
In this section, this paper first introduces channel-wise thresholding for FFN layers. Then, this paper presents the selective sparsification for attention layers. Finally, this paper shows the efficient implementation of the proposed custom sparse kernels.
@@ 3.1 Channel-Wise Thresholding
As described in Equation 4, whether to prune an activation element is determined by both $a^{\text{up}}$ and $a^{\text{gate}}$. To further quantify the significance of each activation element in this decision-making process, we introduce the importance score based on Equation 4,
@@
To obtain all importance scores of elements in Equation 5, we need to compute two matrix-multiplication operations for $a^{\text{gate}}$ and $a^{\text{up}}$. However, the computation for $a^{\text{up}}$ can be reduced by leveraging the sparsity of $a^{\text{gate}}$. Therefore, we need to calculate the score only with $a^{\text{gate}}$. As is shown in Figure 1, we observe that, for each channel $i$, the values of $|a^{\text{up}}_{i}|$ remain relatively consistent across different inputs. However, these values can vary significantly between different channels. Based on this observation, this paper estimates the $|a^{\text{up}}_{i}|$ using the expectation of $|a^{\text{up}}_{i}|$ over sampled training data,
@@
where $n$ is the number of sampled data. Therefore, the importance score is further estimated as,
@@
This threshold indicates the maximal activation value that should be pruned as zero. Different from CATS, this is a Channel-Wise Thresholding (CWT) technique that relates the model performance degradation with the activation sparsity via introducing the importance score in Equation 5.
@@
and the final output of the FFN layer is computed as,
@@ 3.2 Selective Sparsification
where $\mathbf{g}$ and $\mathbf{H}$ denote the first-order and second-order derivatives of the error E with respect to $\hat{x}$, respectively.
@@
Based on Equation 18, this paper also adopts a similar distribution sampling strategy as that in CATS (Lee et al., 2024) to determine the thresholds given a sparsity level. Different from CWT, CATS is a tensor-wise thresholding,
$\text{CATS}(x_{i})=\begin{cases}x_{i},\quad&\text{if}\ |x_{i}|>t\\ 0,\quad&\text{if}\ |x_{i}|\leq t\end{cases}$ (19)
However, which modules the CATS should be applied to becomes a challenge in terms of the trade-off between model performance and model efficiency. The search space is quite large. Taking Llama-7B as an example, which has 32 layers and four attention projections per layer, the search space is over the septillion level.
@@
where $C(\cdot)_{t}$ is the CATS function with the threshold $t$.
Conversely, selective sparsification refers to applying the CATS function to only query and output projections, while not altering key and query projections. The formal expression is,
@@ 3.3 Efficient Sparse Kernels
0:  The sparse input vector $x\in\mathbb{R}^{1\times K}$, the weight matrix $W\in\mathbb{W}^{K\times N}$, the number of output elements $N$, the number of input elements $K$, the block size $B$.
@@
To achieve wall-clock speedup and reduce inference latency based on sparse activations, this paper developed two custom CPU kernels: spvmm (sparse vector-matrix multiplication) and vmmsp (vector-matrix multiplication with output sparsity). The spvmm kernel is optimized for cases where the input activation tensor is sparse, and it is employed in attention modules and FFN down projections. Conversely, the vmmsp kernel is designed for cases where the output activation tensor is multiplied with a sparse mask, and it is used in FFN up projections.
Algorithm 1 and Algorithm 2 show the detailed steps of spvmm and vmmsp, respectively. Algorithm 1 splits the input vector into blocks of size $B$ and accumulates the vector-matrix multiplication results of each block when $x[k]$ is not 0 (Lines 5-7). Algorithm 2 also performs block-level vector-matrix multiplications but computes the outputs at the specific position based on the sparsity mask (Lines 5-9). Both algorithms reduce the latency by bypassing unnecessary weight reads and computations.
The implementation of the vmmsp kernel is relatively straightforward; it computes $Y=XW^{T}$, consistent with the definition of linear projection in PyTorch (Paszke et al., 2019). However, the spvmm kenrel requires a more complex approach to ensure efficient computation on multi-core CPUs while avoiding atomic operations. To this end, we employ two advanced optimizations. First, we employ loop tiling and loop reordering strategies to make sure that each threads compute independently without the need for synchronization or atomic operations. Furthermore, we transpose the linear projection weights in advance during the model preprocessing stage, to memory access continuity and enhance cache hit rates.

### v2

@@ Abstract
Deploying large language models (LLMs) on edge devices presents significant challenges due to the substantial computational overhead and memory requirements. Activation sparsification can mitigate these resource challenges by reducing the number of activated neurons during inference. Existing methods typically employ thresholding-based sparsification based on the statistics of activation tensors. However, they do not model the impact of activation sparsification on performance, resulting in suboptimal performance degradation. To address the limitations, this paper reformulates the activation sparsification problem to explicitly capture the relationship between activation sparsity and model performance. Then, this paper proposes CHESS , a general activation sparsification approach via CHannel-wise thrEsholding and Selective Sparsification. First, channel-wise thresholding assigns a unique threshold to each activation channel in the feed-forward network (FFN) layers. Then, selective sparsification involves applying thresholding-based activation sparsification to specific layers within the attention modules. Finally, we detail the implementation of sparse kernels to accelerate LLM inference. Experimental results demonstrate that the proposed CHESS achieves lower performance degradation over eight downstream tasks while activating fewer parameters than existing methods, thus speeding up the LLM inference by up to 1.27x.
@@
Junhui He1,2,  Shangyu Wu3,4,  Weidong Wen2,  Chun Jason Xue3,  Qingan Li2 thanks: Corresponding author 1 Key Laboratory of Aerospace Information Security and Trusted Computing, Ministry of Education, School of Cyber Science and Engineering, Wuhan University, 2 School of Computer Science, Wuhan University, 3 MBZUAI, 4 City University of Hong Kong
@@ 1 Introduction
Large Language Models (LLMs) have become integral to diverse applications, including code generation, office assistance, voice assistance, and assistive applications for individuals with disabilities. However, due to the substantial computation and memory requirements of LLM inferences, deploying LLMs on edge devices is still challenging. To mitigate these overheads, utilizing the inherent activation sparsity of LLM has emerged as a promising strategy (Liu et al., 2023; Song et al., 2023; Alizadeh et al., 2023). This approach has proven effective for models with the ReLU activation function (Li et al., 2023; Liu et al., 2023).
Contemporary LLMs demonstrate that SwiGLU or GeGLU activation functions can further boost the model performance, but they induce less activation sparsity. Consequently, several methods (Mirzadeh et al., 2024; Song et al., 2024) are proposed to increase sparsity by applying regularization to the activation function and employing continuous training. However, those works require fine-tuning the LLMs, which entails significant overheads. To avoid these overheads and increase activation sparsity in modern LLMs, Lee et al. (2024) propose a thresholding-based pruning method to actively sparsify the activation tensors during the inference stage. However, this thresholding technique focuses solely on the statistics of the activation tensors themselves, failing to model the impact of sparsification on overall model performance, which results in suboptimal performance degradation.
To address the above limitations, this paper proposes CHESS , a new activation sparsification optimization via CHannel-wise thrEsholding and Selective Sparsification. First, this paper reformulates the activation sparsification problem to explicitly capture the relationship between activation sparsity and model performance. Then, this paper proposes channel-wise thresholding for FFN modules in LLMs, which determines a unique threshold for each activation channel. Furthermore, this paper proposes selective sparsification, which applies thresholding-based activation sparsification to the target layers in the attention module. Finally, this paper presents the implementations of sparse kernels to accelerate the inference based on the sparse activations.
To validate the effectiveness of the proposed CHESS , this paper conducts comprehensive experiments on various downstream tasks and state-of-the-art LLMs. Experimental results demonstrate that CHESS can achieve lower performance degradation with better end-to-end inference speedup. The code is publicly available  [^1] https://github.com/ZeonfaiHo/CHESS.
@@
  • This paper reformulates the activation sparsification problem and establishes a connection between sparsity and model performance.
  • This paper proposes two activation sparsification methods, the channel-wise thresholding for FFN modules and the selective sparsification for Attention modules, which can be widely applied in existing LLMs.
  • To make full use of the activation sparsity, this paper details the algorithm design and implementations of the efficient sparse kernels.
  • Experimental results demonstrate the performance and efficiency of the proposed CHESS .
@@ 2.1 Activation Sparsification
Activation functions introduce non-linearity into neural networks, allowing networks to capture complex patterns in the data. ReLU (Glorot et al., 2011), as a popular activation function, has been widely applied in most neural networks to address gradient vanish issues (Zhang et al., 2022). Another benefit of ReLU is introducing the sparsity into the activation tensors. Recent studies (Li et al., 2023; Liu et al., 2023) have demonstrated that up to 95% of the intermediate FFN activations in OPT models are zero. Such sparsity can be used to accelerate the model inference while maintaining comparable model performance (Liu et al., 2023; Alizadeh et al., 2023; Song et al., 2023).
@@
To increase activation sparsity in modern LLMs, existing work (Lee et al., 2024) proposes a thresholding-based pruning method called CATS on some activation tensors in FFN layers. CATS first computes the cutoff threshold over a set of sample input data according to the given sparsity level, then sparsifies the activations during inference and achieves end-to-end speedup via efficient sparse kernel design. Although CATS increases activation sparsity, it only focuses on the statistics of the activation tensors without modeling the impact of activation sparsification on the model performance, leading to suboptimal performance drop.
@@ 2.2 Motivation
Following the observations in CATS (Lee et al., 2024), this paper also aims to apply activation sparsification in the gated-MLP blocks of FFN modules. The formal expression of the FFN module is defined as,
@@
Since the activation function introduces sparsity where the values of many elements in the output tensor are close to zero, we focus on pruning the output of the gate projection layer, i.e., $a^{\text{gate}}$. Then, the following computations, such as the matrix multiplication for $a^{\text{up}}$, the element-wise multiplication between $a^{\text{up}}$ and $a^{\text{gate}}$, or the matrix multiplication with $W^{\text{down}}$, can further be optimized due to the zero elements in the pruned $a^{\text{gate}}$.
@@
We decompose all activations in the pruned tensor into two subsets, i.e., the pruned $\hat{a}^{\text{gate}}_{\mathcal{P}}$ which are all zeros and the non-pruned $\hat{a}^{\text{gate}}_{\mathbb{U}-\mathcal{P}}$ which keep the original values in $a^{\text{gate}}$. Thus, we can simplify the objective defined in Equation 3 as: finding a subset of indices $\mathcal{P}$ that indicates the index of the pruned elements, and satisfies sparsity level $|\mathcal{P}|\geq k\cdot|\mathbb{U}|$, while minimizing the sparsification error illustrated in Equation 4, where $\mathbb{U}=\{1,\ldots,d\}$, $d$ is the feature dimension of $a^{\text{gate}}$.
@@ 3 CHESS : Activation Sparsification via Channel-Wise Thresholding and Selective Sparsification
Figure 1: Distribution of absolute activation values $|a^{\text{up}}_{i}|$ across different inputs for various channels in the FFN of layer 16 of the Llama-3-8B model.
@@
Figure 2: Distribution of $\|W_{i,:}\|_{2}^{2}$ of different rows $i$ in attention projections of layer 16 of Llama-3-8B.
In this section, this paper first introduces the channel-wise thresholding method for FFN modules. Then, it presents the selective sparsification for attention modules. Finally, it discusses the algorithm design and implementations of the efficient custom sparse kernels.
@@ 3.1 Channel-Wise Thresholding
As described in Equation 4, whether to prune an activation element is determined by both $a^{\text{up}}$ and $a^{\text{gate}}$. To quantify the significance of each activation element, we introduce the importance score based on Equation 4,
@@
To obtain all importance scores of elements in Equation 5, we need to compute two matrix-multiplication for $a^{\text{gate}}$ and $a^{\text{up}}$. However, we can reduce the computational cost for $a^{\text{up}}$ by leveraging the sparsity of $a^{\text{gate}}$. Therefore, we need to calculate the score only with $a^{\text{gate}}$. We observe that, for each channel $i$, the values of $|a^{\text{up}}_{i}|$ remain relatively consistent across different inputs, as shown in Figure 1. However, these values can vary significantly between different channels. Based on this observation, this paper estimates the $|a^{\text{up}}_{i}|$ using the expectation of $|a^{\text{up}}_{i}|$ over the sampled input data,
@@
where $n$ is the number of sampled input data. Therefore, the importance score is further estimated as,
@@
This threshold indicates the maximal activation magnitude that should be pruned as zero. Different from CATS, this is a Channel-Wise Thresholding (CWT) technique that relates the model performance with the activation sparsity via introducing the importance score in Equation 5.
@@
and the output of the FFN modules is computed as,
@@ 3.2 Selective Sparsification
where $\mathbf{g}$ and $\mathbf{H}$ denote the first-order and second-order derivatives of the error E with respect to $\hat{x}$, respectively,
@@
Based on Equation 18, this paper adopts the tensor-wise thresholiding method proposed by CATS (Lee et al., 2024) for the projection layers of attention modules:
$\text{CATS}(x_{i})=\begin{cases}0,\quad&\text{if}\ |x_{i}|\leq t\\ x_{i},\quad&\text{if}\ |x_{i}|>t\end{cases}$ (19)
However, which layers the CATS should be applied to becomes a challenge in terms of the trade-off between model performance and model efficiency. The search space is quite large. Taking Llama-2-7B as an example, which has 32 layers and four attention projections per layer, the search space is over the septillion level.
@@
where $C(\cdot)_{t}$ is the CATS function with the threshold $t$. Conversely, selective sparsification refers to applying the CATS function to only query and output projections, while not altering key and query projections. The formal expression is,
@@ 3.3 Efficient Sparse Kernels
0:  The sparse input vector $x\in\mathbb{R}^{1\times K}$, the weight matrix $W\in\mathbb{R}^{K\times N}$, the number of output elements $N$, the number of input elements $K$, the block size $B$.
@@
To achieve wall-clock speedup and reduce inference latency via sparse activations, we developed two custom CPU kernels: spvmm (sparse vector-matrix multiplication) and vmmsp (vector-matrix multiplication with output sparsity). The spvmm kernel is optimized for cases where the input activation tensor is sparse, and it is employed in attention modules and FFN down projections. Conversely, the vmmsp kernel is designed for cases where the output activation tensor is multiplied with a sparse mask, and is used in FFN up projections.
Algorithm 1 and Algorithm 2 show the detailed steps of spvmm and vmmsp, respectively. Both algorithms splits the output vector into blocks of size $B$ and accumulates the results of each block in parallel (Line 1 in Algorithm 1, Line 1 in Algorithm 2). What’s more, they both reduce the latency by bypassing unnecessary weight reads and computations (Line 3 in Algorithm 1, Line 4 in Algorithm 2). Notably, although the implementation of the vmmsp kernel is relatively straightforward, the spvmm kernel requires a more complex approach because its access to each column of $W$ is not continuous. To address this, we employ two advanced optimizations. First, we transpose the linear projection weights in advance during the model preprocessing stage, to ensure memory access continuity. Additionally, we employ loop tiling and loop reordering to make sure that each threads compute independently without the need for synchronization or atomic operations.
4 Experiments
In this section, this paper first introduces the dataset, comparisons, and implementation details. Then, this paper presents the main results over 8 downstream tasks in terms of the model performance and model efficiency. Besides, this paper also conducts an ablation study across different sparsification methods for the attention module and analysis on performance and efficiency over different sparsity level. Additionally, this paper conducts extended comparisons with other state-of-the-art training-free pruning methods, to validate the effectiveness of the proposed CHESS .
4.1 Datasets and Experimental Setup
Datasets We utilize ARC Challenge (Arc-C), ARC Easy (Arc-E), BoolQ, HellaSwag (HS), OpenbookQA (QA), PIQA, SCI-Q, Winogrande (WG) as benchmarks for downstream tasks, employing the Evaluation Harness library from Eleuther AI to ensure consistency with Lee et al. (2024). These tasks are designed to assess various aspects of the language model’s performance, including comprehension, common sense, and reasoning abilities, which effectively illustrate the model’s capability loss with activation sparsification.
Comparisons To validate the effectiveness of the proposed CHESS , we conducted experiments using several state-of-the-art LLMs, including Llama-2-7B, Llama-2-13B, Llama-2-70B, Llama-3-8B and Mistral-7B. These models feature different attention mechanisms, specifically MHA and GQA, and utilize SwiGLU as the activation function for the FFN modules. We tested four different configurations across all five LLMs:
  • Base Model: the LLM model without any activation sparsification.
  • CHESS w/: the channel-wise thresholding and selective sparsification method.
For the ablation study, we evaluate the following three models:
  • Llama-3: the Llama-3 8B model without activation sparsification.
Implementation Details For all models involving activation sparsification, thresholds are sampled from a subset of the C4 dataset (Raffel et al., 2020). Following the settings in CATS (Lee et al., 2024), the sparsity level $k$ is set to 0.5, where the accuracy drop is minimal while the inference latency significantly decreases. The proposed method was implemented using the PyTorch v2.2.2 (Paszke et al., 2019) and HuggingFace Transformers v4.39.3 (Wolf et al., 2019). End-to-end decoding speedups are measured on a randomly collected subset of C4 dataset. Kernel efficiency and end-to-end speedup experiments are conducted with FP32 precision on a personal computer equipped with an Intel Core I9-12900K CPU and 64GB of DDR4 memory. Since our work can be applied to quantized models as well, changing weight precision to FP16 or even lower bit-width quantizations does not materially affect our results (Lee et al., 2024).